Dear brahmsers, We have been trying lately to check out the status of data/calibrations that are so far done for the last run of Au+Au data at 200 AGeV. Truls and Catalin will try to help with the data reduction. From about run 9350 all tracking calibrations are done. For the vertex nothing seems to be done, maybe some run here and there... The same goes for the hodoscopes. Rich seems ok. What about c4? there seems to the a lot of AdcGain and Gap calibrations, but no pedestals... What is the current status??? Furthermore, our colleague David from Gissen will start filling the calibrations for TOF1 and TOF2 in the missing ranges, and Catalin will do this for the BBC, which has to be done first. So, we would like to ask you if you are currently making any of these calibrations? Then we can perhaps divide the work. Thank you! Regards, Truls and Catalin. _______________________________________________ Brahms-dev-l mailing list Brahms-dev-l@lists.bnl.gov http://lists.bnl.gov/mailman/listinfo/brahms-dev-lReceived on Tue Sep 7 11:48:47 2004
This archive was generated by hypermail 2.1.8 : Tue Sep 07 2004 - 11:49:09 EDT